How do I find the radius of a sphere with a volume of 17 cubic feet.
I did V= 4/3 * 3.14 * (8.5) 17 divide by 2 is 8.5 ^3 but I am not getting the right anwer the answer is 1.6 I dnt know how I got it?

Answer by mananth(16949)

V= 17 cu ft
V= 1.33 * pi* r^3
r^3=17/(1.33*pi)
r^3=4.07
r= cuberoot(4.07)
Radius =1.60 ft